# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ 3 The Galaxy Sample, Data Acquisition and Analysis
The galaxies chosen for this program have (1) well-observed SNe Ia with well-defined light curves and peak magnitudes, (2) previously-discovered Cepheid variables (Freedman et al., 2001; Riess et al., 2020), and (3) are located at distances 23 Mpc, suitable for measuring accurate Cepheid, TRGB and JAGB (carbon-star) distances. All three of these methods have high precision and can be independently used to calibrate SNe Ia. Images of our 11 target galaxies are shown in Figure 1.
@@
Observations were obtained using the JWST Near-Infrared Camera (NIRCam; Rieke et al. (2023)), with the filter (comparable to ground-based band) and the F356W filter at 3.6m. The observations were carried out over a 13-month period of time from November 2022 to January 2024. The first observations were obtained for the galaxies NGC 7250, NGC 4536 and NGC 3972 (Lee et al. (2024a); Freedman & Madore (2023a, b); Owens+ (2024a, submitted); Hoyt et al. (2024), making use of the F444W filter at 4.4 m, to both provide a metallicity test for Cepheids, and a long color baseline for the discovery and color discrimination of the JAGB stars. We then switched to F356W for the rest of the target sample, owing to its higher sensitivity and better spatial sampling. However, the F444W filter contains a CO bandhead that is sensitive to metallicity (Scowcroft et al., 2016b), and it is being used to carry out a test for metallicity effects in the two nearest galaxies, M101 and NGC 4258 (at distances 7.5 Mpc) where the highest resolution can be achieved.
@@
NIRCam covers a total area of 9.7 arcmin2 field with a 44 arcsec gap between two 2.2 × 2.2 arcmin regions. Our target fields were chosen to optimize the inclusion of (1) the largest possible number of known Cepheids in the inner disk, (2) portions of the extended outer disk to detect carbon stars, and (3) with a rotation angle optimized for the detection of halo red giants (H24).
@@
A detailed description of our data reduction procedures can be found in Jang+ (2024, in prep), hereafter J24. Here we provide a brief overview. The images333We obtained the stage 2 _cal images with jwst_1149.pmap from the MAST archive. were processed primarily using the NIRCam module in the software package DOLPHOT, updated for JWST data analysis (Weisz et al., 2023, 2024). We found a significant improvement in the point-spread modeling for the newer version of DOLPHOT444DOLPHOT 2.0 NIRCAM module sources (updated 4 Feb 2024), resulting in considerably better source identification and image subtraction using a ‘warmstart’ mode. In addition we are carrying out a parallel and independent analysis using DAOPHOT (Stetson, 1987) in order to provide a quantitative constraint on photometric errors that might arise due to differences in point-spread-function fitting and sky-subtraction approaches in crowded fields. In the case of one of our outer fields in the anchor galaxy, NGC 4258, upon which our current calibration is based, we find excellent agreement between the DOLPHOT and DAOPHOT reductions at a level of 0.002 mag (0.1%) with a total of 0.02 mag down to 3 mag below the TRGB. The comparisons for our other galaxies are ongoing. We also tested the automatic aperture corrections (apcorr = 1) provided by DOLPHOT by comparing the overlapping regions of dithered images, as well as by carrying out comparisons with manual/visual determinations of the aperture corrections. These tests confirmed that the automatic aperture corrections were good to a level of 0.02 mag. Finally, all of our initial data analysis was carried out using a blinding procedure, as described in §4 below.
@@
We note that for our Cepheid analysis, we made use of archival F350LP, F555W and F814W HST data in combination with the F115W JWST near-infrared data. For these archival HST F350LP, F555W, and F814W data, we have adopted the same photometric calibration for the Cepheid fields as given by R22. We did not make use of the HST F160W filter, as the near-infrared resolution with NIRCam JWST surpasses that of HST, and the new data are far less affected by crowding (see Figure 2). The three wavelengths (F555W, F814W and F115W) jointly allow a more precise measurement of the extinction. The Cepheid data were reduced using the ‘warmstart’ mode in DOLPHOT where the x,y coordinates of the sources were determined from a stacked F115W-band image. The optical data were then reduced separately using the same source coordinates (see O24 for more details). For the TRGB and JAGB analyses, we used the NIRCam F115W and the F356W and/or F444W data, and did not use any optical HST data. The source coordinates were obtained from the F115W images, and again used in ‘warmstart’ mode for the longer wavelength data (for more details see H24 and L24).
@@
As described in O24, we injected 1,000 artificial stars into the frames, surrounding each of the known Cepheids. Magnitudes of the artificial stars were retrieved in the same manner as for the real stars in the frames. These measurements then provide a mean (statistical) crowding correction. Application of, and the uncertainties in, these statistical corrections are described in O24.Artificial star corrections were not needed for the TRGB or JAGB measurements. The crowding corrections are most significant for the inner disk (Cepheid) fields. As expected, they are significantly less for JWST than for HST. For the NIRCam filter, the average random (statistical) corrections measured from the artificial stars ranged from -0.003 mag (M101) to -0.012 mag (NGC 3972). For the HST filter, the average corrections ranged from -0.010 mag (NGC 4424) to -0.096 mag (NGC 2442); and for the HST filter, the average corrections ranged from -0.033 mag (M101) to -0.133 mag (NGC 2442).
Galaxy | SN Name | Morphological TypeaaAs described in the text, when combining the methods for the final analysis, the uncertainty for the JAGB method was increased by a factor of 1.18 (given by the ratio of the average values for TRGB and Cepheids to that of the lower value for the JAGB), so that the JAGB method does not unduly contribute more weight than the other two methods. | O/H] bbfootnotemark: | |
@@ 4 Our Blinding Procedure
To set up the blinding procedure initially, the frames were processed using DOLPHOT, as described above. Random numbers were then generated and added separately to each of the photometry catalogs based on the raw data frames. The same random offset was added to all passbands in each galaxy, preserving (arbitrary) colors. The blinding was such that during the entire year and a half of the photometric analysis, no one in the group had any knowledge of what the true distances or the value of might be.
@@
In the case of the Cepheid distance scale, and before unblinding, the R22 Cepheids were identified in the frames, light curves were generated, mean magnitudes and PL relations were measured, artificial star experiments were carried out, and mean bias corrections were applied to the PL relations. The artificial star tests were completed for half the sample before the unblinding step. These tests, which can take several weeks per galaxy, continued until completion, and the corrections were added to the unblinded photometry. The corrections are simply additive, involve no choices, and they did not interfere with the blinding approach.
@@
No further updates were made to the JAGB analysis following unblinding. We can consider the JAGB distances obtained in a completely blinded fashion, and this rung of the blinding experiment a success from start to finish. In the case of the TRGB, all aspects of the analysis remained unchanged after unblinding, except for the sample-wide error estimation and the spatial selection for four of the thirteen fields, all of which had the four largest uncertainties in the blinded analysis. Closer inspection of the color-magnitude diagrams in all of these cases revealed an underpopulated tip that was filled out with the post-unblinding larger spatial selection. (For further details see H24.) The sense of the changes was to decrease the distances in these cases, with a subsequent increase to the value of . These differences were within one-sigma of the blinded results. Finally, subsequent to our unblinding meeting, an error was discovered in our Cepheid analysis pipeline, which resulted in a decrease to our preliminary Cepheid distances (and also a subsequent increase in the value of ). These differences resulted in larger changes at a three-sigma level.
@@ 5 NGC 4258 and Photometric Zero-point Calibration
Our zero-point calibration for the JWST data is based on the geometric distance to the galaxy, NGC 4258. For ground and HST distance scale measurements, the Large Magellanic Cloud (LMC) and the Milky Way also serve as calibrators. Unfortunately, the Milky Way Cepheids, TRGB and JAGB stars are too bright to be observed with JWST. NGC 4258 is a nearby, highly-inclined spiral galaxy located at a distance of 7.6 Mpc. A geometric distance to the galaxy has been measured via its H2O masers, which are orbiting within an accretion disk inclined at 72∘, surrounding a supermassive black hole (see Humphreys et al., 2013; Reid et al., 2019) and megamaser. The most recent geometric distance to NGC 4258 is 29.397 0.024 (stat) 0.022 [sys] mag (Reid et al., 2019), a 1.5% measurement.
@@
An external check of the NGC 4258 calibration can be obtained by comparing previous results for the case where the determination of was based on several anchors (NGC 4258, Milky Way, LMC, and SMC) rather than NGC 4258 alone. In the case of the TRGB, applying distances for these four anchors resulted in (internally) good agreement: Freedman (2021) found values of 69.7 1.0 (stat) 2.0 [sys] for NGC 4258, 69.3 0.8 (stat) 3.5 [sys] for the Milky Way, 69.9 0.5 (stat) 1.6 [sys] for the LMC and 69.5 1.0 (stat) 1.7 [sys] for the SMC calibrations, respectively. Similarly, good internal agreement for the Cepheid calibration was also found by R22 with values of 72.51 1.54 for NGC 4258, 73.02 1.19 for the Milky Way, and 73.59 1.36 for the LMC calibrations, respectively. Thus, having NGC 4258 as the sole calibrator cannot explain the reason for the difference in between 69 and 73: the internal agreement amongst the anchors is excellent for both the TRGB and Cepheids.
@@ 6 Measurement of Distances
In this section we give a brief overview of the three methods used in this study, and compare their advantages and disadvantages. There is no single perfect method. The weaknesses of each method serve to underscore the importance of having independent methods for constraining overall systematic effects in the local distance scale.
@@ 6.1 TRGB
The use of bright red giant branch stars as distance indicators goes back over a century to Harlow Shapley, who measured the brightest stars in globular clusters, as one of his techniques for determining the size of the Milky Way galaxy (Shapley, 1918). Beginning in the 1980s, the availability of CCD detectors with red sensitivity and high quantum efficiency resulted in a resurgence of interest in these stars, this time in their application to the extragalactic distance scale (Mould & Kristian, 1986; Da Costa & Armandroff, 1990; Lee et al., 1993).
@@
The TRGB now provides one of the most precise and accurate means of measuring distances to nearby galaxies, comparable to the Cepheid Leavitt law. In practice, the observed color-magnitude diagrams of old red giant branch stars display a distinct edge/discontinuity in the red giant branch (RGB) luminosity function (LF), corresponding to the core helium-flash luminosity at the end phase of RGB evolution for low-mass stars. Measurement of the TRGB edge has repeatedly been shown to be an excellent standard candle in the I band (Lee et al., 1993; Rizzi et al., 2007; Salaris et al., 2002; Madore et al., 2009; Freedman et al., 2019; Jang et al., 2021), and it is a standardizable candle in the near infrared (Dalcanton et al. (2012); Wu et al. (2014); Madore et al. (2018); Durbin et al. (2020)). Details of the method can be found in a number of reviews (e.g., Madore & Freedman, 1999; Freedman & Madore, 2010; Beaton et al., 2019; Madore & Freedman, 2023).
@@
A strength of the TRGB method is that the underlying theory for why it is an excellent (empirical) standard candle is well understood (Salaris & Cassisi, 1997, 2005; Bildsten et al., 2012; Kippenhahn et al., 2013; Serenelli et al., 2017). After leaving the main sequence, low-mass stars with masses develop a degenerate helium core. They ascend the red giant branch, with their luminosity being powered by a hydrogen-burning shell surrounding the core. The freshly formed helium adds to the core mass until it reaches a value of about 0.5 M⊙ independent of the initial mass of the star. When the temperature reaches about 108K in the isothermal core, it enables the triple-alpha process (helium burning) to commence. Because the degenerate core cannot expand, a thermonuclear runaway ensues (the core helium flash), injecting energy that eventually lifts the electron degeneracy. The star then rapidly evolves away from the red giant branch and descends in luminosity onto the horizontal branch or the red clump, where it undergoes stable core helium burning at a lower-luminosity.
@@
The advantages of the TRGB method are: (1) The discontinuity in the observed TRGB LF is empirically simple to identify and measure. (2) The physical mechanism for the TRGB (the core helium flash) is well understood. (3) If applied in the outer halos of galaxies, where the surface brightness of the galaxy is low, then the overlapping (i.e., crowding or blending effects) of the stellar point spread functions is minimal. (4) In the halos of galaxies, the effects of dust are small (e.g., Ménard et al., 2010). (5) The metallicity of a star on the red giant branch relates directly to its observed color (e.g., Da Costa & Armandroff, 1990; Carretta & Bragaglia, 1998). (6) For a measured TRGB slope in a given passband, the slope of the giant branch luminosity in a different passband is uniquely defined mathematically (Madore & Freedman, 2020).
@@
The disadvantages of the TRGB method are: (1) Care must be taken to ensure that the regions in the target and calibrating galaxies are comparable in terms of surface brightness and line-of-sight column density of dust. (2) Target field placement is critical. Too close in and crowding/blending/reddening become issues. Too far out and the number density of stars is too low for a precise measurement of the tip. (3) Occasionally adjacent peaks of comparable strength can be found in poorly sampled luminosity functions, complicating (or making impossible) an accurate measurement of the true tip. Cases (2) and (3) can be remedied with the acquisition of augmented sample sizes.

### v2

@@ 3 The Galaxy Sample, Data Acquisition and Analysis
The galaxies chosen for this program have (1) well-observed SNe Ia with well-defined light curves and peak magnitudes, (2) previously-discovered Cepheid variables (Freedman et al., 2001; Riess et al., 2022), and (3) are located at distances 23 Mpc, suitable for measuring accurate Cepheid, TRGB and JAGB (carbon-star) distances. All three of these methods are of high precision and can be independently used to calibrate SNe Ia. Images of our 11 target galaxies are shown in Figure 2.
@@
Observations were obtained using the JWST Near-Infrared Camera (NIRCam; Rieke et al. (2023)), with the filter (comparable to ground-based band) and the F356W filter at 3.6m. The observations were carried out over a 13-month period of time from November 2022 to January 2024. The first observations were obtained for the galaxies NGC 7250 and NGC 4536 (Lee et al. (2024a); Freedman & Madore (2023a, b); Hoyt et al. (2024); Owens et al. (2025a)), making use of the F444W filter at 4.4 m, to both provide a metallicity test for Cepheids, and a long color baseline for the discovery and color discrimination of the JAGB stars. We then switched to F356W for the rest of the target sample, owing to its higher sensitivity and better spatial sampling. However, the F444W filter contains a CO bandhead that is sensitive to metallicity (Scowcroft et al., 2016b), and it is being used to carry out a test for metallicity effects in the two nearest galaxies, M101 and NGC 4258 (at distances 7.5 Mpc) where the highest resolution can be achieved.
@@
NIRCam covers a total area of 9.7 arcmin2 field with a 44 arcsec gap between two 2.2 × 2.2 arcmin regions. Our target fields were chosen to optimize the inclusion of (1) the largest possible number of known Cepheids in the inner disk, (2) portions of the extended outer disk to detect carbon stars, and (3) with a rotation angle optimized for the detection of halo red giants (Hoyt et al., 2024).
@@
A detailed description of our data reduction procedures can be found in Jang et al. (2025), hereafter J25. Here we provide a brief overview. The images333We obtained the stage 2 _cal images with jwst_1149.pmap from the MAST archive. were processed primarily using the NIRCam module in the software package DOLPHOT, updated for JWST data analysis (Weisz et al., 2023, 2024). We found a significant improvement in the point-spread modeling for this updated version of DOLPHOT444DOLPHOT 2.0 NIRCAM module sources (updated 4 Feb 2024), resulting in considerably better source identification and image subtraction using a so-called ‘warmstart’ mode. In addition, we are carrying out a parallel and independent analysis of the frames using DAOPHOT (Stetson, 1987) in order to provide a quantitative constraint on photometric errors that might arise due to differences in point-spread-function fitting and sky-subtraction approaches in crowded fields. In the case of one of our outer fields in the anchor galaxy, NGC 4258, upon which our current calibration is based, we find excellent agreement between the DOLPHOT and DAOPHOT reductions at a level of 0.002 mag (0.1%) with a total of 0.02 mag down to 3 mag below the TRGB (Hoyt et al., 2025). The comparisons for our other galaxies are ongoing. We also tested the automatic aperture corrections (Apcor = 1) provided by DOLPHOT by comparing the overlapping regions of dithered images, as well as by carrying out comparisons with manual/visual determinations of the aperture corrections. These tests confirmed that the automatic aperture corrections were good to a level of 0.02 mag. Finally, all of our initial data analysis was carried out using a blinding procedure, as described in §4 below.
@@
We note that for our ongoing Cepheid analysis, we are making use of archival F350LP, F555W and F814W HST data in combination with the F115W JWST near-infrared data. The three wavelengths (F555W, F814W and F115W) jointly allow a more precise measurement of the extinction. However, the optical data (F555W and F814W) are more heavily affected by crowding (see Figure 3). In contrast, for the TRGB and JAGB analyses presented in this paper, we used the NIRCam F115W and the F356W and/or F444W data, and did not use any optical HST data. The source coordinates were obtained from the F115W images, and used in ‘warmstart’ mode for the longer wavelength data (for more details see H25, L25 and J25).
@@
| Galaxy | SN Name | Morphological TypeaaTRGB distances were obtained for four galaxies (NGC 1448, NGC 1559, NGC 5584 and NGC 5643) from F090W archival data discussed in §7.2 (not shown as columns). | O/H] bbfootnotemark: | |
@@ 4 Our Blinding Procedure
To set up the blinding procedure, the frames were processed using DOLPHOT, as described above. Random numbers were then generated and added separately to each of the photometry catalogs based on the raw data frames. The same random offset was added to all passbands in each galaxy, preserving the colors. The blinding was such that during the entire year and a half of the photometric analysis, no one in the group had any knowledge of what the true distances or the value of might be.
@@
In the case of the Cepheid distance scale, subsequent tests after unblinding revealed an issue in the data pipeline and in the photometric reductions of the HST data. A re-analysis of these data is on-going.
@@
No further updates were made to the JAGB analysis following unblinding. We can consider the JAGB distances obtained in a completely blinded fashion, and this rung of the blinding experiment a success from start to finish. In the case of the TRGB, almost all aspects of the analysis remained unchanged after unblinding, except for the sample-wide error estimation and the spatial selection for four of the thirteen fields, all of which had the four largest uncertainties in the blinded analysis. Closer inspection of the color-magnitude diagrams in all of these cases revealed an underpopulated tip that was filled out with the post-unblinding larger spatial selection. (For further details see H25.) The sense of the changes was to decrease the distances in these cases, with a subsequent increase to the value of . These differences were within one-sigma of the blinded results.
@@ 5 NGC 4258 and Photometric Zero-point Calibration
Our zero-point calibration for the JWST data is based on the geometric distance to the galaxy, NGC 4258. For ground and HST distance scale measurements, the LMC and the Milky Way also serve as calibrators. Unfortunately, the Milky Way Cepheids, TRGB and JAGB stars are too bright to be observed with JWST. NGC 4258 is a nearby, highly-inclined spiral galaxy located at a distance of 7.6 Mpc. A geometric distance to the galaxy has been measured via its H2O masers, which are orbiting within an accretion disk inclined at 72∘, surrounding a supermassive black hole (see Humphreys et al., 2013; Reid et al., 2019) and megamaser. The most recent geometric distance to NGC 4258 is 29.397 0.024 (stat) 0.022 [sys] mag (Reid et al., 2019), a 1.5% measurement.
@@
An external check of the NGC 4258 calibration can be obtained by comparing previous results for the case where the determination of was based on several anchors (NGC 4258, Milky Way, LMC, and SMC) rather than NGC 4258 alone. In the case of the TRGB, applying distances for these four anchors resulted in (internally) good agreement: Freedman (2021) found values of 69.7 1.0 (stat) 2.0 [sys] for NGC 4258, 69.3 0.8 (stat) 3.5 [sys] for the Milky Way, 69.9 0.5 (stat) 1.6 [sys] for the LMC and 69.5 1.0 (stat) 1.7 [sys] for the SMC calibrations, respectively. In Appendix A, we undertake a preliminary investigation of the calibration of the JAGB method based on the Milky Way, LMC & SMC, and find very good agreement in the zero-point calibration with that for NGC 4258. Similarly, good internal agreement for the Cepheid calibration was also found by R22 with values of 72.51 1.54 for NGC 4258, 73.02 1.19 for the Milky Way, and 73.59 1.36 for the LMC calibrations, respectively. Thus, having NGC 4258 as the sole calibrator cannot explain the reason for the difference in between 69 and 73: the internal agreement amongst the anchors is excellent for all of the methods.
@@ 6 Measurement of Distances
In this section we give a brief overview of the two methods used in this paper, and discuss their advantages and disadvantages. There is no single perfect method. The weaknesses of each method serve to underscore the importance of having independent methods for constraining overall systematic effects in the local distance scale.
@@ 6.1 TRGB
The use of red giant branch stars as distance indicators goes back over a century to Harlow Shapley, who measured the brightest stars in globular clusters, as one of his techniques for determining the size of the Milky Way galaxy (Shapley, 1918). Beginning in the 1980s, the availability of CCD detectors, with red sensitivity and high quantum efficiency, resulted in a resurgence of interest in these stars, this time in their application to the extragalactic distance scale (Mould & Kristian, 1986; Da Costa & Armandroff, 1990; Lee et al., 1993).
@@
The TRGB now provides one of the most precise and accurate means of measuring distances to nearby galaxies, comparable to the Cepheid Leavitt law. In practice, the observed color-magnitude diagrams of old red giant branch stars display a distinct edge/discontinuity in the red giant branch (RGB) luminosity function (LF). This astrophysical discontinuity corresponds to the core helium-flash luminosity at the end phase of RGB evolution for low-mass stars. Measurement of the TRGB edge has repeatedly been shown to be an excellent standard candle in the band (Lee et al., 1993; Rizzi et al., 2007; Salaris et al., 2002; Madore et al., 2009; Freedman et al., 2019; Jang et al., 2021), and it is a standardizable candle in the near infrared (Dalcanton et al. (2012); Wu et al. (2014); Madore et al. (2018); Durbin et al. (2020); Hoyt et al. (2024)). Details of the method can be found in a number of reviews (e.g., Madore & Freedman, 1999; Freedman & Madore, 2010; Beaton et al., 2019; Madore & Freedman, 2023).
A strength of the TRGB method is that the underlying theory, for why it is an excellent (empirical) standard candle, is well understood (Salaris & Cassisi, 1997, 2005; Bildsten et al., 2012; Kippenhahn et al., 2013; Serenelli et al., 2017; Saltas & Tognelli, 2022). Using a grid of 3 million TRGB models covering a wide range of metallicities and initial masses, Saltas & Tognelli (2022) recently examined the effects of chemical composition, opacity, nuclear reactions, and convection; and compared these results to other published models. Based on a Monte Carlo analysis, they found a maximum theoretical uncertainty of only 1.6%, which is primarily due to radiative opacity systematics. The TRGB is currently the only method for distance determination with a well-understood theoretical basis.
@@
After leaving the main sequence, low-mass stars with masses develop a degenerate helium core. They ascend the red giant branch, with their luminosity being powered by a hydrogen-burning shell surrounding the core. The freshly formed helium adds to the core mass until it reaches a value of about 0.5 M⊙, independent of the initial mass of the star. When the temperature reaches about 108K in the isothermal core, it enables the triple-alpha process (helium burning) to commence. Because the degenerate core cannot expand, a thermonuclear runaway ensues (the core helium flash), injecting energy that eventually lifts the electron degeneracy. The star then rapidly evolves away from the red giant branch and descends in luminosity onto the horizontal branch or the red clump, where it undergoes stable core helium burning at a lower-luminosity.
@@
The advantages of the TRGB method are: (1) The discontinuity in the observed TRGB luminosity function is empirically simple to identify and measure. (2) The physical mechanism for the TRGB (the core helium flash) is well understood. (3) If applied in the outer regions of galaxies, where the surface brightness of the galaxy is low, then the overlapping (i.e., crowding or blending effects) of the stellar point spread functions is minimal. (4) In the outskirts of galaxies, the effects of dust are small (e.g., Ménard et al., 2010). (5) The metallicity of a star on the red giant branch relates directly to its observed color (e.g., Da Costa & Armandroff, 1990; Carretta & Bragaglia, 1998), and can be easily calibrated and corrected for, if required. (6) For a measured TRGB slope in a given passband, the slope of the giant branch luminosity in a different passband is uniquely defined mathematically (Madore & Freedman, 2020).
@@
The disadvantages of the TRGB method are: (1) Care must be taken to ensure that the regions in the target and calibrating galaxies are comparable in terms of surface brightness and line-of-sight column density of dust. (2) Target field placement is critical. Too close in and crowding/blending/reddening become issues. Too far out and the number density of stars is too low for a precise measurement of the tip. (3) Occasionally adjacent peaks of comparable strength can be found in poorly sampled or heavily contaminated luminosity functions, complicating (or making impossible) an accurate measurement of the true tip. Cases (2) and (3) can be remedied with the acquisition of augmented sample sizes.